# THOR2: Leveraging Topological Soft Clustering of Color Space for Human-Inspired Object Recognition in Unseen Environments

### v1

@@ Abstract
Visual object recognition in unseen and cluttered indoor environments is a challenging problem for mobile robots. This study presents a 3D shape and color-based descriptor, TOPS2, for point clouds generated from RGB-D images and an accompanying recognition framework, THOR2. The TOPS2 descriptor embodies object unity, a human cognition mechanism, by retaining the slicing-based topological representation of 3D shape from the TOPS descriptor [1] while capturing object color information through slicing-based color embeddings computed using a network of coarse color regions. These color regions, analogous to the MacAdam ellipses identified in human color perception, are obtained using the Mapper algorithm, a topological soft-clustering technique. THOR2, trained using synthetic data, demonstrates markedly improved recognition accuracy compared to THOR, its 3D shape-based predecessor, on two benchmark real-world datasets: the OCID dataset capturing cluttered scenes from different viewpoints and the UW-IS Occluded dataset reflecting different environmental conditions and degrees of object occlusion recorded using commodity hardware. THOR2 also outperforms baseline deep learning networks, and a widely-used ViT adapted for RGB-D inputs on both the datasets. Therefore, THOR2 is a promising step toward achieving robust recognition in low-cost robots.
@@ 1 Introduction
Object recognition plays a crucial role in robot visual perception, as it supports a wide range of tasks that industrial and domestic use robots must perform, such as manipulation of objects. Object recognition in humans is remarkably advanced; people can identify a wide array of objects in diverse settings, overcoming occlusion-related challenges or changes in appearance, viewpoint, size, scale, or pose. As a step toward achieving comparable performance in robot vision systems with commodity hardware, our previous work [1] presents a 3D shape-based recognition framework, THOR, inspired by a human cognition mechanism known as object unity [2]. In this article, we extend our work to propose the THOR2 framework, which incorporates object color information. Analogous to the MacAdam ellipses [3] that represent colors indistinguishable to an average human, in a first-of-its-kind attempt, we obtain coarse color regions that are considered equivalent to the robot. We obtain these regions using a topological soft-clustering technique known as the Mapper algorithm [4] and use them to compute slicing-based object color embeddings. These embeddings are interleaved with our previously proposed 3D shape-based descriptor to obtain a 3D shape and color-based descriptor.
@@
The primary challenges in developing an object recognition method for low-cost robots used in everyday scenarios involve maintaining consistent performance regardless of varying environmental conditions (such as illumination and background), sensor quality, and the level of clutter in the surroundings. Early deep learning models [5, 6, 7] achieve impressive performance in domain-specific tasks but experience performance degradation in unseen domains with different environmental conditions [8]. Domain adaptation and generalization-based variants of such models have been explored. However, domain adaptation methods require label-agnostic data from the target domain for model training, and domain generalization methods require abundant real-world training data, which poses barriers to deploying such methods on robots with commodity hardware [9]. Therefore, we consider a different paradigm; we consider a single synthetic source domain to obtain descriptors suitable for object recognition across multiple real-world target domains.
@@
Our previous work [1] proposes a topological descriptor computed from depth images, known as TOPS (i.e., Topological features Of Point cloud Slices). Persistent homology is applied to slicing-style filtrations of simplicial complexes constructed from the object’s point cloud to obtain the descriptor. Our approach ensures similarities between the descriptors of the occluded and the corresponding unoccluded objects, embodying object unity. When used with the human-inspired recognition framework, THOR (i.e., TOPS for Human-inspired Object Recognition), the TOPS descriptor shows promising robustness to occlusions. However, recognition using 3D shape information alone is challenging. Additional information cues, such as the color of an object, help achieve improved recognition [10, 11], and are particularly crucial to distinguish between objects with similar geometry.
@@ 5.2.2 THOR2
Training and Testing of Classifiers We use workstations with GeForce GTX 1080 and 1080 Ti GPUs running Ubuntu 18.04 LTS for training (and testing) the classifier models. Separate TOPS-based and TOPS2-based classifiers are trained for every sequence using synthetic RGB-D images in the case of the OCID dataset. In the case of the UW-IS Occluded dataset, we train a single TOPS-based classifier and a single TOPS2-based classifier. For training, first, we augment the training set by mirroring all the view-normalized point clouds across the $x$ and $y$ coordinate axes (in place). The classifiers are five-layer fully connected networks, i.e., Multi-layer Perceptrons (MLPs). During training. we use the Adam optimizer to optimize the categorical cross-entropy loss over 100 epochs. The learning rate is set to $10^{-2}$ for the first 50 epochs and is decreased to $10^{-3}$ for the subsequent 50 epochs.
For testing on the OCID dataset, we use the temporally smoothed, point-wise labeled point clouds provided for every frame in the sequence. In the case of the UW-IS Occluded dataset, which is recorded using commodity hardware and consists of noisy depth images, we consistently perform pre-processing, via voxel grid-based downsampling and outlier removal for every object point cloud, as described in [1]. We report the test results from five-fold cross-validation for THOR2 (and all the comparison methods).
@@ 5.2.3 Comparison methods
Surface normal representations of the depth images for RCFusion, ViT-CNN, and RGB-D ViT are obtained according to [15]. During the training of ViT-CNN and RGB-D ViT, we use the mean and standard deviation corresponding to the ImageNet dataset to standardize the normalized RGB images and surface normal representations of the synthetic dataset. For testing both the models, first, depth interpolation is performed [15], and object crops are obtained using the instance segmentation maps. Subsequently, we replace the background (i.e., background pixels of the RGB image and the corresponding values of the surface normal representation of the depth image) to match that of the synthetic images used for training. The object crops are then resized and padded appropriately to achieve the required input size (i.e., $224\times 224$). Note that we perform the same procedure for background replacement, resizing, and padding RGB-D images from the YCB dataset in the experiments where those real-world images are used to train the RGB-D ViT model.
@@ 5.3.1 Comparison with End-to-end Methods
Evaluation on the UW-IS Occluded dataset Tables 3, 4, and 5 show that THOR2 outperforms all the methods regardless of the environmental condition. Figure 7 and Figure 8 show sample results for the different types of objects under different degrees of occlusion and lighting conditions in the lounge and warehouse environments, respectively.
@@ 5.3.3 Experiments with Real-world Training Data
The previous sections report the recognition performance on real-world data when all the methods are trained using synthetic RGB-D images. By virtue of their topological nature, the TOPS and TOPS2 descriptors used in THOR2 are relatively robust to imprecise depth imagery or illumination-related changes in the real-world data. However, the other end-to-end methods are not explicitly designed to achieve robustness in such issues. Therefore, we perform additional experiments where real-world RGB-D images from the YCB dataset [61] are used (in addition to the synthetic images) to train the best-performing end-to-end model, i.e., RGB-D ViT. Table 7 and Table 8 report the performance of these models on the real-world images of the OCID and the UW-IS Occluded datasets, respectively. We note that THOR2 continues to outperform RGB-D ViT on both the datasets, regardless of the amount of real-world data used for training.
@@
Note: S + xYCB indicates that xreal images from the YCB dataset are used along with the entire synthetic dataset for training and validation.
Table 8: Comparison of mean recognition accuracy over all the object classes (in %) on the UW-IS Occluded dataset under varying degrees of occlusion. THOR2’s performance is compared with RGB-D ViT trained using varying amounts of real-world data in addition to synthetic data.
Environment Occlusion THOR2 RGB-D ViT
Syn only Syn only Syn + 20% YCB Syn + 60% YCB Syn + 100% YCB
Warehouse None 61.40 $\pm$ 0.37 43.65 $\pm$ 0.70 48.14 $\pm$ 1.72 50.17 $\pm$ 1.12 49.39 $\pm$ 2.57
Low 58.00 $\pm$ 0.49 45.11 $\pm$ 1.37 47.56 $\pm$ 1.62 47.38 $\pm$ 1.78 47.25 $\pm$ 3.01
High 59.38 $\pm$ 0.35 42.52 $\pm$ 1.07 48.38 $\pm$ 2.37 47.84 $\pm$ 1.91 46.45 $\pm$ 2.87
Lounge None 64.29 $\pm$ 0.34 39.14 $\pm$ 2.07 44.72 $\pm$ 2.16 46.75 $\pm$ 1.64 46.84 $\pm$ 2.66
Low 65.87 $\pm$ 0.64 43.06 $\pm$ 0.60 47.41 $\pm$ 0.76 47.63 $\pm$ 1.54 47.51 $\pm$ 2.26
High 59.95 $\pm$ 0.52 43.50 $\pm$ 1.07 46.68 $\pm$ 1.75 48.31 $\pm$ 1.78 47.11 $\pm$ 2.90
All 62.58 $\pm$ 0.36 42.96 $\pm$ 0.87 47.04 $\pm$ 1.44 47.97 $\pm$ 1.47 47.41 $\pm$ 2.70
@@
The YCB dataset is recorded using five RGB and depth sensor pairs arranged in a quarter-circular arc. In one experiment, we use 20% of the RGB-D images available for each object to train and validate the RGB-D ViT model. These images are from the first sensor pair placed at table height. In subsequent experiments, we use 60% of the images (i.e., RGB-D images from the first three RGB and depth sensor pairs) and 100% of the images (i.e., RGB-D images from all five RGB and depth sensor pairs placed at e). Table 7 shows that, overall, a small improvement in RGB-D ViT’s performance on the lower camera sequences of the OCID dataset is observed when real-world data is used for training the model. However, in the case of the upper camera sequences, the performance is somewhat lower when 20% of the YCB dataset is used, comparable when the entire YCB dataset is used, and somewhat higher when 60% of the dataset is used. We believe these camera-specific trends can be attributed to the difference in composition of the training set with respect to the camera viewpoints. In the case of the UW-IS Occluded dataset, we observe that RGB-D ViT achieves approximately 4% higher accuracy on the UW-IS Occluded dataset when real-world data is used for training, regardless of the amount of data.
Overall, these results indicate that RGB-D ViT’s performance improves when real-world images are included during training. However, THOR2 continues to outperform RGB-D ViT without using any real-world training data, making it particularly suitable when gathering large volumes of representative real-world training data is infeasible. We refer the reader to Section 6.2 for further discussion of the performance in the context of the sim2real gap between the training and test data.
@@ 6.1 THOR and THOR2 Comparison
The results reported in Section 5.3.1 illustrate that incorporating color information helps THOR2 achieve higher overall performance than THOR on both the OCID and the UW-IS Occluded datasets. However, unlike THOR, THOR2 undergoes a slight drop in performance when the OCID dataset sequences are recorded using the upper camera (see Tables 1 and 2). This observation is consistent with our expectations as THOR’s viewpoint-based classifier library equips it for recognition in such scenarios where challenging object poses are more likely to occur (for e.g., S-08). Despite the drop in performance, THOR2 achieves higher overall accuracy than THOR. Moreover, the fewer classifiers in THOR2 make it more easily extendable to an open-world scenario where new object classes may appear. For instance, when new classes appear, a class-adaptive extension of THOR2 (with two classifiers) would be more efficiently updated than THOR, which requires updating all the classifiers in the library.

### v2

# THOR2: Topological Analysis for 3D Shape and Color-Based Human-Inspired Object Recognition in Unseen Environments

@@ Abstract
Visual object recognition in unseen and cluttered indoor environments is a challenging problem for mobile robots. This study presents a 3D shape and color-based descriptor, TOPS2, for point clouds generated from RGB-D images and an accompanying recognition framework, THOR2. The TOPS2 descriptor embodies object unity, a human cognition mechanism, by retaining the slicing-based topological representation of 3D shape from the TOPS descriptor [1] while capturing object color information through slicing-based color embeddings computed using a network of coarse color regions. These color regions, analogous to the MacAdam ellipses identified in human color perception, are obtained using the Mapper algorithm, a topological soft-clustering technique. THOR2, trained using synthetic data, demonstrates markedly improved recognition accuracy compared to THOR, its 3D shape-based predecessor, on two benchmark real-world datasets: the OCID dataset capturing cluttered scenes from different viewpoints and the UW-IS Occluded dataset reflecting different environmental conditions and degrees of object occlusion recorded using commodity hardware. THOR2 also outperforms baseline deep learning networks, and a widely-used Vision Transformer (ViT) adapted for RGB-D inputs trained using synthetic and limited real-world data on both the datasets. Therefore, THOR2 is a promising step toward achieving robust recognition in low-cost robots.
@@ 1 Introduction
Object recognition plays a crucial role in robot visual perception, as it supports a wide range of tasks that industrial and domestic use robots must perform, such as manipulation of objects. Object recognition in humans is remarkably advanced; people can identify a wide array of objects in diverse settings, overcoming occlusion-related challenges or changes in appearance, viewpoint, size, scale, or pose. People can also distinguish between individual instances of objects, as, for example, a softball from a tennis ball, even though both are balls. As a step toward achieving comparable performance in robots with commodity hardware, our previous work [1] presents a 3D shape-based recognition framework, THOR, inspired by a human cognition mechanism known as object unity [2]. In this article, we extend our work to propose the THOR2 framework, which incorporates object color information. Analogous to the MacAdam ellipses [3] that represent colors indistinguishable to an average human, in a first-of-its-kind attempt, we obtain coarse color regions that are considered equivalent to the robot. We obtain these regions using a topological soft-clustering technique known as the Mapper algorithm [4] and use them to compute slicing-based object color embeddings. These embeddings are interleaved with our previously proposed 3D shape-based descriptor to obtain a 3D shape and color-based descriptor.
@@
The primary challenges in developing an object instance recognition method for low-cost robots used in everyday scenarios involve maintaining consistent performance regardless of varying environmental conditions (such as illumination and background), sensor quality, and the level of clutter in the surroundings. Early deep learning models [5, 6, 7] achieve impressive performance in domain-specific tasks but experience performance degradation in unseen domains with different environmental conditions [8]. Domain adaptation and generalization-based variants of such models have been explored. However, domain adaptation methods require label-agnostic data from the target domain for model training, and domain generalization methods require abundant real-world training data, which poses barriers to deploying such methods on robots with commodity hardware [9]. Therefore, we consider an alternative paradigm: we consider a single synthetic source domain to obtain the descriptors suitable for object instance recognition across multiple real-world target domains.
@@
Our previous work [1] proposes a topological descriptor computed from depth images, known as TOPS (i.e., Topological features Of Point cloud Slices). Persistent homology is applied to slicing-style filtrations of simplicial complexes constructed from the object’s point cloud to obtain the descriptor. Our approach ensures similarities between the descriptors of the occluded and the corresponding unoccluded objects, embodying object unity. When used with the human-inspired recognition framework, THOR (i.e., TOPS for Human-inspired Object Recognition), the TOPS descriptor shows promising robustness to occlusions. However, recognition using 3D shape information alone is challenging. Additional information cues, such as the color of an object, help achieve improved recognition [10, 11], and are particularly crucial to distinguish among the objects instances with similar geometry.
@@ 5.2.2 THOR2
Training and Testing of Classifiers We use workstations with GeForce GTX 1080 and 1080 Ti GPUs running Ubuntu 18.04 LTS for training (and testing) the classifier models. Separate TOPS-based and TOPS2-based classifiers are trained for every sequence using synthetic RGB-D images in the case of the OCID dataset. In the case of the UW-IS Occluded dataset, we train a single TOPS-based classifier and a single TOPS2-based classifier. For training, first, we augment the training set by mirroring all the view-normalized point clouds across the $x$ and $y$ coordinate axes (in place). The classifiers are five-layer fully connected networks, i.e., Multi-layer Perceptrons (MLPs) with layers having 512, 256, 128, 64, and n_classes nodes, respectively, where n_classes represents the number of object classes. During training, we use the Adam optimizer to optimize the categorical cross-entropy loss over 100 epochs. The learning rate is set to $10^{-2}$ for the first 50 epochs and is decreased to $10^{-3}$ for the subsequent 50 epochs.
For testing on the OCID dataset, we use the temporally smoothed, point-wise labeled point clouds provided for every frame in the sequence. The UW-IS Occluded dataset is recorded using commodity hardware and consists of noisy depth images. Instead of existing approaches for denoising depth images [67], we consistently perform pre-processing via voxel grid-based downsampling and outlier removal for every object point cloud, as described in [1]. We report the test results from five-fold cross-validation for THOR2 (and all the comparison methods).
@@ 5.2.3 Comparison methods
Surface normal representations of the depth images for RCFusion, ViT-CNN, and RGB-D ViT are obtained according to [15]. During the training of ViT-CNN and RGB-D ViT, we use the mean and standard deviation corresponding to the ImageNet dataset to standardize the normalized RGB images and surface normal representations of the synthetic dataset. For testing both the models, first, depth interpolation is performed [15], and object crops are obtained using the instance segmentation maps. Subsequently, we replace the background (i.e., background pixels of the RGB image and the corresponding values of the surface normal representation of the depth image) to match that of the synthetic images used for training. The object crops are then resized and padded appropriately to achieve the required input size (i.e., $224\times 224$). Note that we perform the same procedure for background replacement, resizing, and padding RGB-D images from the YCB dataset in the experiments where those real-world images are used to train the RCFusion, ViT-CNN, and RGB-D ViT models.
@@ 5.3.1 Comparison with End-to-end Methods
Evaluation on the UW-IS Occluded dataset Table 3, 4, and 5 show that THOR2 outperforms all the methods regardless of the environmental condition. Figure 7 and Figure 8 show sample results for the different types of objects under different degrees of occlusion and lighting conditions in the lounge and warehouse environments, respectively.
@@ 5.3.3 Experiments with Real-world Training Data
The previous sections report the recognition performance on real-world data when all the methods are trained using synthetic RGB-D images. By virtue of their topological nature, the TOPS and TOPS2 descriptors used in THOR2 are relatively robust to imprecise depth imagery or illumination-related changes in the real-world data. However, the other end-to-end methods are not explicitly designed to achieve robustness in such issues. Therefore, we perform additional experiments where real-world RGB-D images from the YCB dataset [61] and synthetic images are used for training the end-to-end models. Table 7 reports the performance of RGB-D ViT on the OCID dataset, when it is trained using synthetic images and different amounts of real-world images from the YCB dataset. Table 8 reports the performance of all the end-to-end methods, namely, RCFusion, ViT-CNN, and RGB-D ViT, on the UW-IS Occluded dataset, when they are trained using synthetic images and the entire YCB dataset. We note that THOR2, trained entirely on synthetic data, continues to outperform the end-to-end methods on both the OCID and UW-IS Occluded datasets.
@@
Table 8: Comparison of mean recognition accuracy over all the object classes (in %) on the UW-IS Occluded dataset under varying degrees of occlusion. THOR2’s performance is compared with the end-to-end methods, namely, RCFusion, ViT-CNN, and RGB-D ViT trained using real-world data from the YCB dataset and synthetic data.
Environment Occlusion THOR2 RCFusion ViT-CNN RGB-D ViT
Syn only Syn + 100% YCB Syn + 100% YCB Syn + 100% YCB
Warehouse None 61.40 $\pm$ 0.37 36.57 $\pm$ 3.07 14.47 $\pm$ 0.18 49.39 $\pm$ 2.57
Low 58.00 $\pm$ 0.49 34.44 $\pm$ 3.01 13.30 $\pm$ 0.51 47.25 $\pm$ 3.01
High 59.38 $\pm$ 0.35 32.57 $\pm$ 3.15 13.35 $\pm$ 0.37 46.45 $\pm$ 2.87
Lounge None 64.29 $\pm$ 0.34 40.37 $\pm$ 2.89 13.17 $\pm$ 0.45 46.84 $\pm$ 2.66
Low 65.87 $\pm$ 0.64 35.42 $\pm$ 1.95 11.24 $\pm$ 0.41 47.51 $\pm$ 2.26
High 59.95 $\pm$ 0.52 28.00 $\pm$ 2.83 15.41 $\pm$ 0.40 47.11 $\pm$ 2.90
Average accuracy across all scenarios 62.58 $\pm$ 0.36 34.42 $\pm$ 2.66 13.47 $\pm$ 0.36 47.41 $\pm$ 2.70
Increase in accuracy due to YCB data - 13.58 1.02 4.45
The YCB dataset is recorded using five RGB and depth sensor pairs arranged in a quarter-circular arc. In one experiment, we use 20% of the RGB-D images available for each object to train and validate the RGB-D ViT model. These images are from the first sensor pair placed at table height. In subsequent experiments, we use 60% of the images (i.e., RGB-D images from the first three RGB and depth sensor pairs) and 100% of the images (i.e., RGB-D images from all five RGB and depth sensor pairs placed at e). Table 7 shows that, overall, a small improvement in RGB-D ViT’s performance on the lower camera sequences of the OCID dataset is observed when real-world data is used for training the model. However, in the case of the upper camera sequences, the performance is somewhat lower when 20% of the YCB dataset is used, comparable when the entire YCB dataset is used, and somewhat higher when 60% of the dataset is used. We believe these camera-specific trends can be attributed to the difference in composition of the training set with respect to the camera viewpoints.
In the case of the UW-IS Occluded dataset, we observe that all the end-to-end methods achieve higher recognition accuracy when real-world images from the YCB dataset are added to the training data. Overall, RGB-D ViT achieves approximately 4.5% higher accuracy, and the highest increase is when none of the objects in the test images are occluded. This observation is expected given that the training and test data distributions are most alike in this case; both training and test data have real-world images of unoccluded objects. Similar observations can be made for RCFusion, although the gain in accuracy is substantially higher as compared to RGB-D ViT. In the case of ViT-CNN, the gain is minimal as the method does not consider retraining/fine-tuning of the feature extraction stage.
Overall, these results indicate that the performance of end-to-end models improves when real-world images are included during training. However, THOR2 continues to outperform them without using any real-world training data, making it particularly suitable when gathering large volumes of representative real-world training data is infeasible. We refer the reader to Section 6.2 for further discussion in the context of the sim2real gap between the training and test data.
@@ 6.1 THOR and THOR2 Comparison
The results reported in Section 5.3.1 illustrate that incorporating color information helps THOR2 achieve higher overall performance than THOR on both the OCID and the UW-IS Occluded datasets. However, unlike THOR, THOR2 undergoes a slight drop in performance when the OCID dataset sequences are recorded using the upper camera (see Table 1 and 2). This observation is consistent with our expectations as THOR’s viewpoint-based classifier library equips it for recognition in such scenarios where challenging object poses are more likely to occur (for e.g., S-08). Despite the drop in performance, THOR2 achieves higher overall accuracy than THOR. Moreover, the fewer classifiers in THOR2 make it more easily extendable to an open-world scenario where new object classes may appear. For instance, when new classes appear, a class-adaptive extension of THOR2 (with two classifiers) would be more efficiently updated than THOR, which requires updating all the classifiers in the library.
@@ 6.2 Sim2Real Gap
6.3 Separability of Objects
Separability of objects, i.e., how distinct and easily identifiable different objects are within a feature space, plays a key role in determining how challenging it can be to recognize them. Good descriptors can enhance the separability of objects, leading to improved recognition. In the case of the OCID dataset, every sequence considers a different set of objects capturing different separability scenarios. Therefore, Table 1, Table 2, and Table 7 consider each sequence separately and train separate models for each sequence. We also perform experiments where a single model is trained for all the sequences instead of treating each sequence individually. Table 9 reports recognition accuracy from these experiments for sequence recordings obtained using the lower camera. We note that the overall performance for all the methods drops compared to Table 1, where all the sequences are considered separately. Such an outcome is expected, especially in the case of object instance recognition, because of the inherent similarities among the objects in the entire dataset. Overall, THOR2 continues to outperform all the other methods in this case as well.
Table 9: Comparison of mean recognition accuracy (in %) of THOR2 with the end-to-end models on the OCID dataset sequences recorded using the lower camera (best in bold), when all the objects appearing in all the sequences are considered during training.
Place Scene type Seq. ID RCFusion ViTCNN RGB-D ViT THOR THOR2
Table Curved S-25 11.78 $\pm$ 2.76 15.07 $\pm$ 3.45 28.47 $\pm$ 5.28 21.78 $\pm$ 2.37 37.00 $\pm$ 2.06
S-26 6.89 $\pm$ 3.69 9.78 $\pm$ 0.65 43.69 $\pm$ 8.19 32.42 $\pm$ 4.53 48.85 $\pm$ 4.12
S-35 0.00 $\pm$ 0.00 9.54 $\pm$ 1.13 10.93 $\pm$ 3.52 19.10 $\pm$ 1.87 41.98 $\pm$ 3.26
S-36 4.20 $\pm$ 2.33 20.00 $\pm$ 2.04 44.27 $\pm$ 7.44 28.12 $\pm$ 3.31 29.78 $\pm$ 2.69
Cuboid S-23 47.04 $\pm$ 6.09 41.68 $\pm$ 4.19 53.04 $\pm$ 2.02 60.14 $\pm$ 3.47 70.66 $\pm$ 2.45
S-24 44.01 $\pm$ 7.27 33.71 $\pm$ 3.84 46.92 $\pm$ 3.92 54.95 $\pm$ 1.23 73.73 $\pm$ 1.83
S-33 34.17 $\pm$ 5.09 34.17 $\pm$ 3.77 57.57 $\pm$ 5.60 52.65 $\pm$ 2.98 58.99 $\pm$ 4.00
S-34 35.43 $\pm$ 5.30 38.63 $\pm$ 3.97 58.97 $\pm$ 5.52 52.95 $\pm$ 1.22 70.95 $\pm$ 1.77
Mixed S-21 27.67 $\pm$ 4.42 28.75 $\pm$ 2.33 19.47 $\pm$ 2.92 50.90 $\pm$ 3.84 60.82 $\pm$ 3.25
S-22 12.31 $\pm$ 3.00 11.29 $\pm$ 1.76 41.88 $\pm$ 6.88 57.40 $\pm$ 4.81 64.93 $\pm$ 1.91
S-31 30.84 $\pm$ 5.05 32.00 $\pm$ 1.17 60.89 $\pm$ 2.81 55.26 $\pm$ 1.76 70.42 $\pm$ 3.93
S-32 25.01 $\pm$ 3.77 48.47 $\pm$ 2.58 21.31 $\pm$ 6.09 55.94 $\pm$ 4.29 60.29 $\pm$ 2.37
Floor Curved S-05 2.22 $\pm$ 2.22 19.81 $\pm$ 3.99 16.11 $\pm$ 7.11 19.42 $\pm$ 4.21 58.77 $\pm$ 3.04
S-06 14.89 $\pm$ 1.88 30.96 $\pm$ 3.55 21.73 $\pm$ 1.76 39.31 $\pm$ 4.34 65.86 $\pm$ 4.14
S-11 12.44 $\pm$ 0.89 23.52 $\pm$ 1.43 29.39 $\pm$ 1.47 31.06 $\pm$ 5.86 55.72 $\pm$ 5.83
S-12 22.58 $\pm$ 0.81 30.43 $\pm$ 1.47 49.28 $\pm$ 1.94 35.35 $\pm$ 2.33 37.66 $\pm$ 6.95
Cuboid S-03 31.98 $\pm$ 2.55 26.15 $\pm$ 4.34 47.98 $\pm$ 5.54 59.17 $\pm$ 3.65 82.74 $\pm$ 2.53
S-04 31.22 $\pm$ 3.72 18.55 $\pm$ 3.32 40.31 $\pm$ 2.58 59.65 $\pm$ 1.69 77.81 $\pm$ 1.63
S-09 49.60 $\pm$ 2.28 32.00 $\pm$ 2.66 41.49 $\pm$ 1.71 56.10 $\pm$ 6.01 76.67 $\pm$ 2.50
S-10 33.04 $\pm$ 4.73 23.91 $\pm$ 4.14 44.88 $\pm$ 3.93 71.74 $\pm$ 2.38 90.37 $\pm$ 2.95
Mixed S-01 16.04 $\pm$ 3.56 39.00 $\pm$ 1.99 27.84 $\pm$ 4.08 57.23 $\pm$ 4.35 86.23 $\pm$ 3.74
S-02 20.19 $\pm$ 4.36 25.69 $\pm$ 0.77 37.70 $\pm$ 0.58 58.26 $\pm$ 3.40 52.56 $\pm$ 3.90
S-07 37.98 $\pm$ 7.42 45.98 $\pm$ 1.45 60.87 $\pm$ 3.27 56.21 $\pm$ 0.53 62.16 $\pm$ 1.93
S-08 21.93 $\pm$ 1.92 25.11 $\pm$ 3.04 44.31 $\pm$ 3.14 16.19 $\pm$ 3.15 34.74 $\pm$ 3.89
All sequences 29.17 $\pm$ 1.00 28.75 $\pm$ 0.35 40.03 $\pm$ 0.91 46.90 $\pm$ 0.68 59.96 $\pm$ 0.63